OpenStudy (anonymous):

write an equation of the line containing the given points and parallel to the given line (4,7); x+7y=6

OpenStudy (anonymous):

Once again, find the slope of the given line which is -1/7 Can you see why? Since in parallel, the slopes are equal , the slope of the line which we are going to write also has the same slope -1/7. Can u proceed from now on..........

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\] Now put m=-1/7 ans x1=4 and y1=7 in the given formula

OpenStudy (anonymous):

Okay one sec let me try to work this

OpenStudy (anonymous):

ugh I suck working at this :(